OpenStudy (howard-wolowitz):

converge or diverge? If the sequence does converge, what number do the terms of the sequence converge to?

OpenStudy (howard-wolowitz):

The sequence does not converge. The sequence converges to 100. The sequence converges to -100. The sequence converges to zero.

OpenStudy (bunnykitten02):

cool name

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (anonymous):

What's the value of \[\lim_{n\to\infty}(-1)^n\frac{100n}{100+n}\,?\]

OpenStudy (howard-wolowitz):

100n

OpenStudy (anonymous):

No, you can find the limit by comparing the coefficients of the \(n\) terms in the numerator and denominator. Meanwhile, \(\lim_{n\to\infty}(-1)^n=\pm1\). \[\lim_{n\to\infty}(-1)^n\frac{100n}{100+n}=\left(\lim_{n\to\infty}(-1)^n\right)\left(\lim_{n\to\infty}\frac{100n}{100+n}\right)=\pm\lim_{n\to\infty}\frac{100n}{100+n}\]

OpenStudy (howard-wolowitz):

so you think its B

OpenStudy (howard-wolowitz):

Because it has 100 in it

OpenStudy (howard-wolowitz):

ohhh ok so u can find the limit

OpenStudy (howard-wolowitz):

so it does converges

OpenStudy (anonymous):

Sure, \(100\) is the limit of the second expression, but that doesn't mean the sequence converges to \(100\). We still have that \(\pm\) sign to worry about. This means the sequence alternates between \(-100\) and \(100\) as \(n\) grows and grows. The fact that \(a_n\) switches between two values forever like this means the sequence cannot converge.

OpenStudy (howard-wolowitz):

well how do we get ride of that stupid sign?

OpenStudy (howard-wolowitz):

i think its a postive / negative sign right or infinity

OpenStudy (anonymous):

We don't! It tells us exactly what we want to know about this sequence: It does not converge.

OpenStudy (howard-wolowitz):

oh oko wel thcxs for expaning that for me

OpenStudy (howard-wolowitz):

I'm gooing with A then

OpenStudy (anonymous):

yw

OpenStudy (howard-wolowitz):

ready for next question?